OpenStudy (anonymous):

Let f(x) = 3x+2 and g(x) = x^2 Part 1 Find f(g(x)) Part 2 Find g(f(x))

OpenStudy (anonymous):

f(x) = 4x+2 g(x) = x^2 f(g(x)) = f(x^2) g(f(x)) = g( 3x+2)

OpenStudy (anonymous):

f(g(x)) f(g(x)) = 3(x^2) + 2 = 3x^2 + 2 g(f(x)) = (3x+2)^2 = 9x^2 + 12x + 4